Factor the polynomial and use the factored form to find the zeros. (Enter your answers as a comma-separated list. Enter all answers including repetitions.)
P(x) = x3 + 3x2 − 9x − 27

Answer by lwsshak3(11628)

Factor the polynomial and use the factored form to find the zeros. (Enter your answers as a comma-separated list. Enter all answers including repetitions.)
P(x) = x3 + 3x2 − 9x − 27
P(x) =x^2(x+3)-9(x+3)
P(x)=(x+3)(x^2-9)